OpenStudy (anonymous):

Four cups of a salad blend containing 40% spinach is mixed with an unknown amount of a salad blend containing 55% spinach. The resulting salad contains 50% spinach. How many cups of salad are in the resulting mixture? 8 9 12 13

OpenStudy (anonymous):

@shinalcantara

OpenStudy (confluxepic):

Yes.

OpenStudy (shinalcantara):

let 'x' be the amount @ 40% 'y' be the amount @ 55% then you'll have 0.40x + 0.55y = 0.50(x+y) -----equation 1 0.40(4) + 0.55y = (0.50)(4+y) get the value of y

OpenStudy (shinalcantara):

remember that the resulting mixture amounts to x+y

OpenStudy (anonymous):

8 ?

OpenStudy (shinalcantara):

yes and the the resulting mixture is?

OpenStudy (anonymous):

8

OpenStudy (shinalcantara):

x + y ?

OpenStudy (shinalcantara):

x being 4 and y=8 resulting mixture = x + y = 4 + 8 =12

OpenStudy (anonymous):

Oooh .

OpenStudy (shinalcantara):

in dealing with mixture problems, all you have to do is to identify the amount and the concentration and the resulting mixture

OpenStudy (shinalcantara):

say in this problem, you have 4cups @ 40% concentration, y-cups @ 55% concentration. 0.40(4) + 0.55y. since the desired mixture is 50% concentrated then you can say that adding the amounts 4cups and y-cups would result to 50% mixture then you'll have it as 0.40(4) + 0.55y = 0.50(4 + y)